OpenStudy (jojokiw3):

Question in comments. How do I do this Algebraic question with imaginary number "i"? Steps please.

OpenStudy (jojokiw3):

\[\frac{ \sqrt{-16}i^{3}-4i^{2} }{ 1-\sqrt{-4} }\]

OpenStudy (anonymous):

ok i might be a little rusty on this but heres what i got

OpenStudy (anonymous):